OpenStudy (anonymous):

Integral xsqrt(9-16x^2) dx Integral xsqrt(9-16x^2) dx @MIT 18.01 Single …

OpenStudy (lalaly):

let u=9-16x^2 du=-32xdx \[\large{\frac{-1}{32} \int\limits{\sqrt{u}du} = \frac{-1}{32} (\frac{2}{3}) u^{\frac{3}{2}}}\]\[=\huge{\frac{-1}{48}} (9-16x^2)^{\frac{3}{2}}\]

OpenStudy (anonymous):

i know how to do this, i was giving this as an example for someone

OpenStudy (lalaly):

i know, i was trying to do it since no one tried